# Quantum Algorithms for the Pathwise Lasso

### v2

@@ 2.2 Concentration bounds
Fact 8 (Hoeffding’s bound).
@@
Fact 9 (Lévy’s lemma, [3, Corollary 4.4.28]).
@@ 2.4 Quantum subroutines
Fact 13 (Quantum minimum finding [25]).
@@
Fact 15 ([10, Theorem 12]).
Given $M\in\mathbb{N}$ and access to an $(n+1)$-qubit unitary $U$ satisfying
$U|0^{n}\rangle|0\rangle=\sqrt{a}|\psi_{1}\rangle|1\rangle+\sqrt{1-a}|\psi_{0}\rangle|0\rangle,$
where $|\psi_{0}\rangle$ and $|\psi_{1}\rangle$ are arbitrary $n$-qubit states and $a\in(0,1)$, there is a quantum algorithm that uses $O(M)$ applications of $U$ and $U^{\dagger}$ and $\widetilde{O}(M)$ elementary gates, and outputs a state $|\Lambda\rangle$ such that, after measuring $|\Lambda\rangle$, with probability at least $9/10$, the first register $\lambda$ of the outcome satisfies
@@
Finally, the following result, which is based on [23, Theorem 3.4], constructs the mapping $|j\rangle|\bar{0}\rangle\mapsto|j\rangle|R_{j}\rangle$ such that $R_{j}$ holds an approximation to $\mathbf{A}_{j}^{\top}\mathbf{u}$ up to an additive error, where $\mathbf{A}\in\mathbb{R}^{n\times d}$ and $\mathbf{u}\in\mathbb{R}^{n}$ can be accessed quantumly via KP-trees.
@@
Fix $j\in[d]$. We start by approximating $\mathbf{A}_{j}^{\top}\mathbf{u}=\sum_{i=1}^{n}A_{ij}u_{i}$ up to error $\epsilon\|\mathbf{A}_{j}\|_{\infty}\|\mathbf{u}\|_{1}$. Apply the operator from Fact 11 to create
@@
Here, $|{\phi_{1}}\rangle$ and $|{\phi_{0}}\rangle$ are unit vectors. We now use Fact 15 to get an estimate $\widetilde{a}_{+}$ such that
@@
Sample $q=\lceil 2\epsilon^{-2}\ln(2/\delta)\rceil$ indices $\{i_{1},\dots,i_{q}\}\subseteq[n]$ from the distribution $\mathcal{D}_{\mathbf{A}_{j}}$ by using sampling access to $\mathbf{A}$ and set $z_{k}=\|\mathbf{A}_{j}\|_{1}u_{i_{k}}\operatorname{sign}(A_{i_{k}j})$, $k\in[q]$. The algorithm outputs $\widehat{z}=\frac{1}{q}\sum_{k=1}^{q}z_{k}$ as an estimate for $\mathbf{A}_{j}^{\top}\mathbf{u}$. Since $z_{1},\dots,z_{q}$ are i.i.d. copies of $z$, a Hoeffding’s bound (Fact 8) gives
@@ 3.1 Lasso solution
In this section, we review several known properties of the solution to the Lasso regression problem defined as
$\hat{\bm{\beta}}\in\operatorname*{argmin}_{\bm{\beta}\in\mathbb{R}^{d}}\frac{1}{2}\|\mathbf{y}-\mathbf{X}{\bm{\beta}}\|_{2}^{2}+\lambda\|\bm{\beta}\|_{1}\quad\text{for some}\leavevmode\nobreak\ \lambda>0,$
where $\mathbf{X}\in\mathbb{R}^{n\times d}$ is the design matrix and $\mathbf{y}\in\mathbb{R}^{n}$ is a vector of observations. We will cover optimality conditions for the Lasso problem, the general form for its solution, and continuity and uniqueness conditions. Most of these results can be found in [91, 78, 104, 92]. We start with obtaining the optimality conditions for the Lasso problem.
Fact 18 ([104, Lemma 1]).
@@
Equation 5 can be obtained by considering the Karush–Kuhn–Tucker conditions, or sub-gradient optimality conditions. More specifically, a vector $\hat{\bm{\beta}}\in\mathbb{R}^{d}$ is a solution to the Lasso problem if and only if (see, e.g., [12, Proposition 3.1.5])
$\mathbf{0}\in\{\mathbf{X}^{\top}(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}})-\lambda\mathbf{s}:\mathbf{s}\in\partial\|\hat{\bm{\beta}}\|_{1}\},$
@@
Moving on, first note that $\bm{\eta}$ is a sub-gradient and that the uniqueness of $\mathbf{X}\hat{\bm{\beta}}$ (Fact 18) implies the uniqueness of $\mathcal{A}$ and $\bm{\eta}$. By the definition of the sub-gradients $\mathbf{s}$ in Equation 5, we know that $\hat{\bm{\beta}}_{\mathcal{I}}=\mathbf{0}$, and therefore the block $\mathcal{A}$ of Equation 5 can be written as
@@
By Fact 19, any Lasso solution has $\ell_{2}$-norm
@@
It follows from Fact 19 that the Lasso solution is unique if $\operatorname{null}(\mathbf{X}_{\mathcal{A}})=\{\mathbf{0}\}$ and is given by
@@
It is known [104, Lemma 14] that there always is a Lasso solution such that $|\mathcal{A}|\leq\min\{n,d\}$. If the Lasso solution is unique, then the active set has size at most $\min\{n,d\}$. The sufficient condition $\operatorname{null}(\mathbf{X}_{\mathcal{A}})=\{\mathbf{0}\}$ for uniqueness has appeared several times in the literature [79, 39, 113, 18, 104]. Since the active set $\mathcal{A}$ depends on the Lasso solution at $\mathbf{y},\mathbf{X},\lambda$, the condition that $\mathbf{X}_{\mathcal{A}}$ has full rank is somewhat circular. Because of this, more natural conditions are assumed which imply $\operatorname{null}(\mathbf{X}_{\mathcal{A}})=\{\mathbf{0}\}$, e.g., the general position property [88, 28, 30, 104] and the matrix $\mathbf{X}$ being drawn from a continuous probability distribution [104]. We say that a matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ has columns in general position if no $k$-dimensional subspace in $\mathbb{R}^{n}$, for $k<\min\{n,d\}$, contains more than $k+1$ elements from the set {±$\{\pm\mathbf{X}_{1},\dots,\pm\mathbf{X}_{d}\}$, excluding antipodal pairs. It is known [88, 28, 30, 104] that this is enough to guarantee the uniqueness of the Lasso solution.
Fact 21 ([104, Lemma 3]).
@@
It is well known that, when $\operatorname{null}(\mathbf{X}_{\mathcal{A}})=\{\mathbf{0}\}$, the KKT conditions from Fact 19 imply that the regularisation path $\mathcal{P}\triangleq\{\hat{\bm{\beta}}(\lambda)\in\mathbb{R}^{d}:\lambda>0\}$ comprising all the solutions for $\lambda>0$ is continuous piecewise linear, which was first proven by Efron, Hastie, Johnstone, and Tibshirani [32], cf. Equation 2. This fact was later extended to the case when $\operatorname{rank}(\mathbf{X}_{\mathcal{A}})<|\mathcal{A}|$ by Tibshirani [104], who proved that the path of solutions with $\mathbf{b}=\mathbf{0}$ in Equation 6 is continuous and piecewise linear.
Fact 23 ([104, Appendix A]).
@@ 3.2 The LARS algorithm
Starting at $\lambda=\infty$, where the Lasso solution is trivially $\mathbf{0}\in\mathbb{R}^{d}$, the LARS algorithm iteratively computes all the kinks in the Lasso solution path by decreasing the parameter $\lambda$ and checking for points where the path’s linear trajectory changes. This is done by making sure that the KKT conditions from Fact 19 are always satisfied. When a kink is reached, a coordinate from the solution $\hat{\bm{\beta}}(\lambda)$ will go from non-zero (being in the active set $\mathcal{A}$) to zero (join the inactive set $\mathcal{I}$), or vice-versa, i.e., from zero (leave $\mathcal{I}$) to non-zero (into $\mathcal{A}$). We shall now describe the LARS algorithm in more details, following [104]. Algorithm 1 summarises the LARS algorithm.
@@
$\lambda_{t+1}^{\rm join}=\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}(\mathcal{A})\},\qquad i_{t+1}^{\rm join}=\operatorname*{argmax}_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}(\mathcal{A})\}.$
@@
$\lambda_{t+1}^{\rm cross}=\max_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}(\mathcal{A})\},\qquad i_{t+1}^{\rm cross}=\operatorname*{argmax}_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}(\mathcal{A})\}.$
@@ 4.1 Simple quantum LARS algorithm
12      $i_{t+1}^{\rm join}\leftarrow\operatorname{argmax}_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}\}$ and $\lambda_{t+1}^{\rm join}\leftarrow\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}\}$ with failure probability $\frac{\delta}{T}$ (Fact 13)
@@ 4.2 Approximate quantum LARS algorithm
If a vector $\widetilde{\bm{\beta}}\in\mathbb{R}^{d}$ satisfies the $\mathsf{KKT}_{\lambda}(\epsilon)$ condition for $\epsilon\geq 0$, then $\widetilde{\bm{\beta}}$ minimises the Lasso cost function up to an error $\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$, i.e.,
@@
is called a duality gap and provides an optimality guarantee (see, e.g., [12, Section 4.3]):
@@
12      Obtain $\widetilde{i}_{t+1}^{\rm join}\in\big{\{}j\in\mathcal{I}:\Lambda_{j}^{\rm join}\geq\big{(}1-\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\big{)}\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}\}\big{\}}$ with failure probability $\frac{\delta}{T}$ (Fact 14 and Lemma 16)
@@ 24      $t\leftarrow t+1$
We now present our approximate quantum LARS Algorithm 3 for the pathwise Lasso. Its main idea is quite simple: we improve the search of the joining variable $i_{t+1}^{\rm join}$ and time $\lambda_{t+1}^{\rm join}$ over $i\in\mathcal{I}$ by approximately computing the joining times $\Lambda_{i}^{\rm join}(\mathcal{A})=\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}$ and using the approximate quantum minimum-finding subroutine from Chen and de Wolf [23] (Fact 14). More specifically, the joining variable $\widetilde{i}_{t+1}^{\rm join}$ is obtained to be any variable $j\in\mathcal{I}$ that maximises $\Lambda_{j}^{\rm join}(\mathcal{A})$ up to some small relative error $\epsilon$, i.e., at any iteration, Algorithm 3 randomly samples from the set
@@
After asserting the correctness of Algorithm 3, we now analyse its complexity. Specifically, we show how to sample the joining variable $\widetilde{i}_{t+1}^{\rm join}$ from the set in Equation 11 by combining the approximate quantum minimum-finding subroutine from Chen and de Wolf [23] (Fact 14) and the unitary map that approximates (parts of) the joining times $\Lambda_{i}^{\rm join}(\mathcal{A})=\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}$ (Lemma 16). Our final complexity depends on the mutual incoherence between different column subspaces and the overlap between $\mathbf{y}$ and different column subspaces.
@@
then $\epsilon_{0}\leq\frac{\epsilon}{2(1+\alpha_{\mathcal{A}})}\lambda^{\rm join}_{t+1}$ in time $O\Big{(}\frac{\gamma_{\mathcal{A}}^{-1}+\sqrt{n}\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}}{(1-\alpha_{\mathcal{A}})\epsilon}\log(1/\delta_{0})\operatorname{poly}\log(nd)\Big{)}$. Finally, we apply the approximate quantum minimum-finding algorithm (Fact 14) to obtain an index $\widetilde{i}_{t+1}^{\rm join}\in\mathcal{I}$ such that

### v3

@@ 2.2 Concentration bounds
Fact 5 (Hoeffding’s bound).
@@
Fact 6 (Lévy’s lemma, [3, Corollary 4.4.28]).
@@ 2.4 Quantum subroutines
Fact 10 (Quantum minimum finding [26]).
@@
Fact 12 ([11, Theorem 12]).
Given $M\in\mathbb{N}$ and access to an $(n+1)$-qubit unitary $\mathbf{U}$ satisfying
$\mathbf{U}|0^{n}\rangle|0\rangle=\sqrt{a}|\psi_{1}\rangle|1\rangle+\sqrt{1-a}|\psi_{0}\rangle|0\rangle,$
where $|\psi_{0}\rangle$ and $|\psi_{1}\rangle$ are arbitrary $n$-qubit states and $a\in(0,1)$, there is a quantum algorithm that uses $O(M)$ applications of $\mathbf{U}$ and $\mathbf{U}^{\dagger}$ and $\widetilde{O}(M)$ elementary gates, and outputs a state $|\Lambda\rangle$ such that, after measuring $|\Lambda\rangle$, with probability at least $9/10$, the first register $\lambda$ of the outcome satisfies
@@
Finally, the following result, which is based on [24, Theorem 3.4], constructs the mapping $|j\rangle|\bar{0}\rangle\mapsto|j\rangle|R_{j}\rangle$ such that $R_{j}$ holds an approximation to $\mathbf{A}_{j}^{\top}\mathbf{u}$ up to an additive error, where $\mathbf{A}\in\mathbb{R}^{n\times d}$ and $\mathbf{u}\in\mathbb{R}^{n}$ can be accessed quantumly via KP-trees.
@@
Fix $j\in[d]$. We start by approximating $\mathbf{A}_{j}^{\top}\mathbf{u}=\sum_{i=1}^{n}A_{ij}u_{i}$ up to error $\epsilon\|\mathbf{A}_{j}\|_{\infty}\|\mathbf{u}\|_{1}$. Apply the operator from Fact 8 to create
@@
Here, $|{\phi_{1}}\rangle$ and $|{\phi_{0}}\rangle$ are unit vectors. We now use Fact 12 to get an estimate $\widetilde{a}_{+}$ such that
@@
Sample $q=\lceil 2\epsilon^{-2}\ln(2/\delta)\rceil$ indices $\{i_{1},\dots,i_{q}\}\subseteq[n]$ from the distribution $\mathcal{D}_{\mathbf{A}_{j}}$ by using sampling access to $\mathbf{A}$ and set $z_{k}=\|\mathbf{A}_{j}\|_{1}u_{i_{k}}\operatorname{sign}(A_{i_{k}j})$, $k\in[q]$. The algorithm outputs $\widehat{z}=\frac{1}{q}\sum_{k=1}^{q}z_{k}$ as an estimate for $\mathbf{A}_{j}^{\top}\mathbf{u}$. Since $z_{1},\dots,z_{q}$ are i.i.d. copies of $z$, a Hoeffding’s bound (Fact 5) gives
@@ 3.1 Lasso solution
In this section, we review several known properties of the solution to the Lasso regression problem with inputs $(\mathbf{X},\mathbf{y})\in\mathbb{R}^{n\times d}\times\mathbb{R}^{n}$ defined as
$\hat{\bm{\beta}}\in\operatorname*{arg\,min}_{\bm{\beta}\in\mathbb{R}^{d}}\mathcal{L}^{(\lambda)}_{\mathbf{X},\mathbf{y}}(\bm{\beta})\quad\text{where}\quad\mathcal{L}^{(\lambda)}_{\mathbf{X},\mathbf{y}}(\bm{\beta})\triangleq\frac{1}{2}\|\mathbf{y}-\mathbf{X}{\bm{\beta}}\|_{2}^{2}+\lambda\|\bm{\beta}\|_{1}\quad\text{for}\leavevmode\nobreak\ \lambda>0.$ (5)
Here $\mathbf{X}\in\mathbb{R}^{n\times d}$ is the design matrix and $\mathbf{y}\in\mathbb{R}^{n}$ is a vector of observations. When the inputs are clear from context, we shall simply write $\mathcal{L}^{(\lambda)}(\bm{\beta})$. We will cover optimality conditions for the Lasso problem, the general form for its solution, and continuity and uniqueness conditions. Most of these results can be found in [93, 80, 106, 94]. We start with obtaining the optimality conditions for the Lasso problem.
Fact 15 ([106, Lemma 1]).
@@
Equation 6 can be obtained by considering the Karush–Kuhn–Tucker conditions, or sub-gradient optimality conditions. More specifically, a vector $\hat{\bm{\beta}}\in\mathbb{R}^{d}$ is a solution to the Lasso problem if and only if (see, e.g., [13, Proposition 3.1.5])
$\mathbf{0}\in\{-\nabla\mathcal{L}^{(\lambda)}(\hat{\bm{\beta}})=\mathbf{X}^{\top}(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}})-\lambda\mathbf{s}:\mathbf{s}\in\partial\|\hat{\bm{\beta}}\|_{1}\},$
@@
Moving on, first note that $\bm{\eta}$ is a sub-gradient and that the uniqueness of $\mathbf{X}\hat{\bm{\beta}}$ (Fact 15) implies the uniqueness of $\mathcal{A}$ and $\bm{\eta}$. By the definition of the sub-gradients $\mathbf{s}$ in Equation 6, we know that $\hat{\bm{\beta}}_{\mathcal{I}}=\mathbf{0}$, and therefore the block $\mathcal{A}$ of Equation 6 can be written as
@@
By Fact 16, any Lasso solution has $\ell_{2}$-norm
@@
It follows from Fact 16 that the Lasso solution is unique if $\operatorname{null}(\mathbf{X}_{\mathcal{A}})=\{\mathbf{0}\}$ and is given by
@@
It is known [106, Lemma 14] that there always is a Lasso solution such that $|\mathcal{A}|\leq\min\{n,d\}$. If the Lasso solution is unique, then the active set has size at most $\min\{n,d\}$. The sufficient condition $\operatorname{null}(\mathbf{X}_{\mathcal{A}})=\{\mathbf{0}\}$ for uniqueness has appeared several times in the literature [81, 40, 115, 19, 106]. Since the active set $\mathcal{A}$ depends on the Lasso solution at $\mathbf{y},\mathbf{X},\lambda$, the condition that $\mathbf{X}_{\mathcal{A}}$ has full rank is somewhat circular. Because of this, more natural conditions are assumed which imply $\operatorname{null}(\mathbf{X}_{\mathcal{A}})=\{\mathbf{0}\}$, e.g., the general position property [90, 29, 31, 106] and the matrix $\mathbf{X}$ being drawn from a continuous probability distribution [106]. We say that a matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ has columns in general position if no $k$-dimensional subspace in $\mathbb{R}^{n}$, for $k<\min\{n,d\}$, contains more than $k+1$ elements from the set {±$\{\pm\mathbf{X}_{1},\dots,\pm\mathbf{X}_{d}\}$, excluding antipodal pairs. It is known [90, 29, 31, 106] that this is enough to guarantee the uniqueness of the Lasso solution.
Fact 18 ([106, Lemma 3]).
@@
It is well known that, when $\operatorname{null}(\mathbf{X}_{\mathcal{A}})=\{\mathbf{0}\}$, the KKT conditions from Fact 16 imply that the regularisation path $\mathcal{P}\triangleq\{\hat{\bm{\beta}}(\lambda)\in\mathbb{R}^{d}:\lambda>0\}$ comprising all the solutions for $\lambda>0$ is continuous piecewise linear, which was first proven by Efron, Hastie, Johnstone, and Tibshirani [33], cf. Equation 2. This fact was later extended to the case when $\operatorname{rank}(\mathbf{X}_{\mathcal{A}})<|\mathcal{A}|$ by Tibshirani [106], who proved that the path of solutions with $\mathbf{b}=\mathbf{0}$ in Equation 7 is continuous and piecewise linear.
Fact 20 ([106, Appendix A]).
@@ 3.2 The LARS algorithm
Starting at $\lambda=\infty$, where the Lasso solution is trivially $\mathbf{0}\in\mathbb{R}^{d}$, the LARS algorithm iteratively computes all the kinks in the Lasso solution path by decreasing the parameter $\lambda$ and checking for points where the path’s linear trajectory changes. This is done by making sure that the KKT conditions from Fact 16 are always satisfied. When a kink is reached, a coordinate from the solution $\hat{\bm{\beta}}(\lambda)$ will go from non-zero (being in the active set $\mathcal{A}$) to zero (join the inactive set $\mathcal{I}$), or vice-versa, i.e., from zero (leave $\mathcal{I}$) to non-zero (into $\mathcal{A}$). We shall now describe the LARS algorithm in more details, following [106]. Algorithm 1 summarises the LARS algorithm.
@@
$\lambda_{t+1}^{\rm join}=\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}(\mathcal{A})\}\quad\text{and}\quad i_{t+1}^{\rm join}=\operatorname*{argmax}_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}(\mathcal{A})\}.$
@@
$\lambda_{t+1}^{\rm cross}=\max_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}(\mathcal{A})\}\quad\text{and}\quad i_{t+1}^{\rm cross}=\operatorname*{argmax}_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}(\mathcal{A})\}.$
@@ 4.1 Simple quantum LARS algorithm
12      $i_{t+1}^{\rm join}\leftarrow\operatorname{argmax}_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}\}$ and $\lambda_{t+1}^{\rm join}\leftarrow\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}\}$ with failure probability $\frac{\delta}{T}$ (Fact 10)
@@ 4.2 Approximate quantum LARS algorithm
If a vector $\widetilde{\bm{\beta}}\in\mathbb{R}^{d}$ satisfies the $\mathsf{KKT}_{\lambda}(\epsilon)$ condition for $\epsilon\geq 0$, then $\widetilde{\bm{\beta}}$ minimises the Lasso cost function $\mathcal{L}^{(\lambda)}(\bm{\beta})$ up to an error $\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$, i.e.,
@@
is called a duality gap and provides an optimality guarantee (see, e.g., [13, Section 4.3]):
@@
Remark 25.
Even though the error $\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$ might look unnatural, it can be relaxed to a relative error by observing that it implies
$\mathcal{L}^{(\lambda)}(\widetilde{\bm{\beta}})-\min_{\bm{\beta}\in\mathbb{R}^{d}}\mathcal{L}^{(\lambda)}(\bm{\beta})\leq\epsilon\lambda\|\widetilde{\bm{\beta}}\|_{1}\leq\epsilon\mathcal{L}^{(\lambda)}(\widetilde{\bm{\beta}})$ $\iff(1-\epsilon)\big{(}\mathcal{L}^{(\lambda)}(\widetilde{\bm{\beta}})-\min_{\bm{\beta}\in\mathbb{R}^{d}}\mathcal{L}^{(\lambda)}(\bm{\beta})\big{)}\leq\epsilon\min_{\bm{\beta}\in\mathbb{R}^{d}}\mathcal{L}^{(\lambda)}(\bm{\beta})$
$\iff\mathcal{L}^{(\lambda)}(\widetilde{\bm{\beta}})-\min_{\bm{\beta}\in\mathbb{R}^{d}}\mathcal{L}^{(\lambda)}(\bm{\beta})\leq\frac{\epsilon}{1-\epsilon}\min_{\bm{\beta}\in\mathbb{R}^{d}}\mathcal{L}^{(\lambda)}(\bm{\beta}).$
We now present our approximate quantum LARS Algorithm 3 for the pathwise Lasso. Its main idea is quite simple: we improve the search of the joining variable $i_{t+1}^{\rm join}$ and time $\lambda_{t+1}^{\rm join}$ over $i\in\mathcal{I}$ by approximately computing the joining times $\Lambda_{i}^{\rm join}(\mathcal{A})=\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}$ and using the approximate quantum minimum-finding subroutine from Chen and de Wolf [24] (Fact 11). More specifically, the joining variable $\widetilde{i}_{t+1}^{\rm join}$ is obtained to be any variable $j\in\mathcal{I}$ that maximises $\Lambda_{j}^{\rm join}(\mathcal{A})$ up to some small relative error $\epsilon$, i.e., at any iteration, Algorithm 3 randomly samples from the set
@@
12      Obtain $\widetilde{i}_{t+1}^{\rm join}\in\big{\{}j\in\mathcal{I}:\Lambda_{j}^{\rm join}\geq\big{(}1-\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\big{)}\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}\}\big{\}}$ with failure probability $\frac{\delta}{T}$ (fact 11 and 13)
@@ 24      $t\leftarrow t+1$
After asserting the correctness of Algorithm 3, we now analyse its complexity. Specifically, we show how to sample the joining variable $\widetilde{i}_{t+1}^{\rm join}$ from the set in Equation 12 by combining the approximate quantum minimum-finding subroutine from Chen and de Wolf [24] (Fact 11) and the unitary map that approximates (parts of) the joining times $\Lambda_{i}^{\rm join}(\mathcal{A})=\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}$ (Lemma 13). Our final complexity depends on the mutual incoherence between different column subspaces and the overlap between $\mathbf{y}$ and different column subspaces.
@@
then $\epsilon_{0}\leq\frac{\epsilon}{2(1+\alpha_{\mathcal{A}})}\lambda^{\rm join}_{t+1}$ in time $O\Big{(}\frac{\gamma_{\mathcal{A}}^{-1}+\sqrt{n}\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}}{(1-\alpha_{\mathcal{A}})\epsilon}\log(1/\delta_{0})\operatorname{poly}\log(nd)\Big{)}$. Finally, we apply the approximate quantum minimum-finding algorithm (Fact 11) to obtain an index $\widetilde{i}_{t+1}^{\rm join}\in\mathcal{I}$ such that